Database last updated on Mon, 02 Feb 2026 10:10:52 +0000 by localhost
Instrument constant changes for the week before 2023308 (1 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 15 rows)
Measurements and tests summary.
| _jday | _loc | _serial | _skc | aode | ds | ed | hg | sl | um | uv | zs |
| 308 | TORONTO | #015 | Next schedule will be to2mk2um (now to2mk2um ZS= 98 ZS= 98),to2mk2um,Waiting until ZA = 98 now 94.7,Waiting until ZA = 98 now 96.4 | 8 | 8 | 2 | 49 | 17 | 80 | 69 | 42 |
(0 sec)
Dead time (1 sec)
Nothing to report.
Standard lamp R6 (1 sec)
| File name | Count | Serial | Expected (+/-20) | Measured | 30-day history |
| B30823.015 | 17 | #015 | 1990 | 1799 |